Calculate the wavelengths in air at 20?C for sounds in the maximum range of human hearing. The speed of sound in air is 343 m/s. What is the wavelength of a 24-MHz ultrasonic wave?
Answer #1

wavelength= v/ f= 343/ 24*10^6= 14.3 *10^-6 m
Answer #2

You would use the same formula as in the first part: wavelength = speed of sound divided by frequency.

So, using 343 m/s for the speed of sound, the wavelength would be:

343 m/s / 24 X 10^6 Hz = 1.429 X 10^-5 meters, or 14.29
Answer #3

speed=wavelength*frequency

343=24*106*wavelength

wavelength=14.29*10-6

=1.429*10-5
Answer #4


wavelength = V/f = 343/(24e6)


= 14.29*10^-6 m